OpenStudy (anonymous):

This is for an algebra 2 exam review and i don't know do you distribute first or use the exponent first?? Problem: 2(x+3)^(2/3) - 8 = 24

OpenStudy (watchmath):

Add 8 to both side \(2(x+3)^{2/3}=32\) Divide both sides by 2 \((x+3)^{2/3}=16\) Raise both side to the 3/2 power \(x+3=64\) \(x=61\)

OpenStudy (anonymous):

Thank you very much!!

OpenStudy (anonymous):

Good method. It took me a minute to figure it out.